Grade 12Physical Chemistry

what is the wave no. for the longest wavelength transition in lyman series of atomic hydrogen spectrum

nidhi jaiswal
8 Years agoGrade 12

1 Answer

Umakant Biswal
8 Years ago
Dear nidhi 
apply the equation 
wave number = Rh ( 1/ n(1) ^2 – 1/(n2)^2 ) 
here u mentioned about lyman series and longest wavelength 
so, n1 = 1 and n2=2 
so, by putting the value 
wave number = 109678 ( 1/(1)^2 – 1/(2)^2 ) 
wave number = 109678  ( 1/1 – 1/4) 
= 109678 ( 1-1/4 ) = 109678 * ¾ 
kindly calculate the above and u will get the answer 
HOPE IT CLEARS YOUR DOUBT 
CHEERS